OpenStudy (anonymous):

16,4,1,1/4 ... Which exponential function describes the given sequence?

OpenStudy (anonymous):

A) y = 1 4 x - 1 B) y = 4x - 1 C) y = 16· 1 4 x - 1 D) y = 16·4x - 1